# Demons registration for 2D empirical wavelet transform: Application to texture segmentation

### v2

@@ 4.3 Mapping estimation assessment
We assess the behavior of the different algorithms to map a partition to a generating function’s Fourier support which is usually a disk or a square. Table 1 reports the average RMSE of the estimated mappings $\widetilde{\gamma}_{n}$ from the sets $\Omega_{n}$ of the Watershed or Voronoi partition to the disk or square $\Lambda$ for the different demons algorithms. The additive and Thirion’s demons have similar average performance, but the additive one suffers from less variability, and both outperform the diffeomorphic demons due to the latter’s additional constraint on diffeomorphicity. In addition, the performance of all demons is better for the more shape-constrained Voronoi partition than for the Watershed partition, and better for the disk than for the square due to its irregularity.
@@
To assess the invertibility, continuity and diffeomorphicity of the mapping estimates, we explore the behavior of the resulting empirical wavelet coefficients expected to be concave and equal to one on most of each support $\Omega_{n}$. Figure 4 and Figure 5 show the disk band-pass empirical wavelet coefficients $\widehat{\chi}_{n}^{\mathrm{D},\tau}$, where $\tau=0.2$, with and without normalization, respectively, for the Watershed partition and the different demons algorithms. The normalized coefficients are not concave for the additive and Thirion’s demons due to the discontinuity of the normalization coefficient while they are concave but very sparse for the diffeomorphic demons. The unnormalized wavelet coefficients are more satisfactory as they are continuous, concave and mostly equal to one for each $\Omega_{n}$. However, the coefficients obtained using the diffeomorphic demons are more spread out due to the additional constraint of differentiability. As for the Thirion’s demons, it can miss an important part of a filter (see the top right corner of Figure 5).
@@
In conclusion, the additive demons provides accurate continuous mapping estimates to compute robust unnormalized empirical wavelet filters in a reasonable computation time. As for the diffeomorphic demons, it provides inaccurate diffeomorphic mapping estimates, leading to unsuitable normalized filters.
@@ 4.4 Reconstruction assessment
To assess the numerical behavior of the empirical wavelet transform, which is theoretically lossless for the disk and square band-pass filter according to Theorem 2, we examinate the quality of reconstruction. Table 3 reports the PSNR of the empirical wavelet reconstruction of the toy image for the disk and square band-pass filters with different transition widths $\tau$, the Watershed and Voronoi partitions and the different demons algorithms. The additive demons provides better or at least similar reconstruction than the other demons. Moreover, the normalization in the empirical wavelet filters has no impact on the reconstruction performance for the different demons.
Despite the accuracy of the mapping estimation obtained in Table 1, the related wavelet reconstruction can be corrupted when the estimated filter bank is not entirely covering the Fourier domain. Notably, the transition width $\tau=0.2$ permits to ensure a larger covering of the Fourier domain than the transition width $\tau=0.1$ by the filters in case of an inaccurate estimate of a mapping. In contrast, the transition width $\tau=0.3$ can lead to artefacts in a symmetric wavelet $\chi_{n}$ due to overlaps of paired wavelet filters $\widehat{\psi}_{n}$ and $\widehat{\psi}_{-n}$, as observed for the disk band-pass transform using the Watershed partition and a diffeomorphic demons.
@@ 4.5 Application to texture segmentation
Analyzing the structure and chemical properties of self-assembled monolayers of organic molecules is the core of many applications in nanoscience and nanotechnology [17, 10]. These properties result in variations in the textures that are crucial to identify. Recently, Empirical Curvelet Wavelet Transform (EWT-Curvelet), obtained by partitioning the Fourier domain in scales and angles [9], has shown to provide relevant texture features for scanning tunneling microscope images of self-assembled monolayers [11, 2]. In this section, we explore the relevance of the proposed Empirical Wavelet Transform for this task.
used the proposed approach to characterize patterns of cyanide molecules [11, 12]
We perform the texture segmentation according to the procedure of [14], which consists of the following steps. First, we extract the textural part of the images by performing a cartoon–texture decomposition [1, 8]. Next, we perform an empirical wavelet transform and compute the local energy of the obtained coeficients which is the average of the coefficients over a window of size $19\times 19$ at each pixel. Finally, segmentation is obtained by performing a pixel-wise k-means clustering with cityblock distance for preselected numbers of cluster $k$ on the local energy.
Figure 6 shows partitions of the Fourier spectrum of a scanning tunneling microscope image obtained by the Curvelet, Watershed and Voronoi methods with a scale-space step-size $s_{0}=0.8$ on the logarithm of the Fourier spectrum. The Curvelet method results in a large number of sets $\Omega_{n}$, namely $393$, while the Watershed and Voronoi methods extract as many sets $\Omega_{n}$ as detected harmonic modes, $17$ in this case.
@@
Figure 7 shows scanning tunneling microscope images, their corresponding texture parts, and their segmentation for the EWT-Curvelet and the proposed disk and square band-pass empirical wavelet transforms with $\tau=0.2$ for both Voronoi and Watershed partitioning. The proposed empirical wavelet transforms provide accurate and comparable texture masks for the different images. In all cases, the segmentation performance with the proposed approach is better than with the EWT-Curvelet, as can be clearly seen in the last two rows of Figure 7.
@@ 5 Discussion
The proposed numerical implementation based on the additive demons algorithm provides accurate continuous mappings suitable for the empirical wavelet transform. The results show that this approach has satisfactory performances for reconstruction and texture extraction. In contrast, diffeomorphic mapping estimates are harder and longer to obtain. This suggests to avoid in practice the use of normalization proposed in previous work [18].
The proposed algorithm applies for any homogeneous or separable wavelet kernel and any Fourier partition. However, the selection of the wavelet kernel’s support and the partitioning method is crucial to optimize the mapping estimation performance. The study conducted in the present work shows that a Voronoi partition and a wavelet kernel defined on a disk result in a better performance overall.
The proposed numerical procedure is ready for application on real-world data. In particular, it has shown to provide relevant texture features for scanning tunneling microscope images for the different studied wavelet kernels. In addition, the number of features resulting from the Voronoi or Watershed partitioning methods is reasonnable compared to the one obtained by the Curvelet transform used in [2].
Future work will include a thorough study of the stop criterion of the additive demons algorithm and the use of different wavelet kernels in applications.
@@ 6 Conclusions
This work proposed an efficient algorithm for the empirical wavelet transform from any wavelet kernel based on demons registration. Among the different meticulously compared demons algorithms, the additive one has shown to provide accurate numerical empirical wavelet systems with good properties of reconstruction, as validated for suitable wavelet kernels. The relevance of this approach for texture feature extraction has also been highlighted. A Matlab toolbox will be made publicly available at the time of publication.

### v3

@@ 4.3 Mapping estimation assessment
We assess the behavior of the different algorithms to map a partition to a generating function’s Fourier support, which is usually a disk or a square. Table 1 reports the average RMSE of the estimated mappings $\widetilde{\gamma}_{n}$ from the sets $\Omega_{n}$ of the Watershed or Voronoi partition to the disk or square $\Lambda$ for the different demons algorithms. The additive and Thirion’s demons algorithms have similar performance on average, and both outperform the diffeomorphic demons algorithm due to the latter’s additional constraint on diffeomorphicity. However, the Thirion’s demons algorithm suffers from more variability than the additive demons algorithm. Furthermore, all demons algorithms achieve higher performance for the more shape-constrained Voronoi partition than for the Watershed partition and better for the disk than for the square due to its irregularity.
@@
To assess the invertibility, continuity, and diffeomorphicity of the mapping estimates, we explore the behavior of the resulting empirical wavelet coefficients expected to be concave and equal to one on most of each support $\Omega_{n}$. Figures 4 and 5 show the disk band-pass empirical wavelet coefficients $\widehat{\chi}_{n}^{\mathrm{D},\tau}$, where $\tau=0.2$, respectively, with and without normalization for the Watershed partition and the different demons algorithms. The normalized coefficients are not concave for the additive and Thirion’s demons algorithms due to the discontinuity of the normalization coefficient, while they are concave but very sparse for the diffeomorphic demons algorithm. The unnormalized wavelet coefficients are more satisfactory as they are continuous, concave and mostly equal to one for all sets $\Omega_{n}$. However, the coefficients obtained using the diffeomorphic demons algorithm are more spread out due to the additional constraint of differentiability. As for the Thirion’s demons algorithm, it can miss an important part of a filter (see the top right corner of Figure 5).
@@
In conclusion, the additive demons algorithm provides accurate continuous mapping estimates to compute robust unnormalized empirical wavelet filters in a reasonable computation time. As for the diffeomorphic demons algorithm, it provides inaccurate diffeomorphic mapping estimates, leading to unsuitable normalized filters.
@@ 4.4 Reconstruction assessment
To assess the numerical behavior of the empirical wavelet transform, which is theoretically lossless for the disk and square band-pass filter according to Theorem 2, we examine the quality of reconstruction. Table 3 reports the PSNR of the empirical wavelet reconstruction of the toy image for the disk and square band-pass filters with different transition widths $\tau$, the Watershed and Voronoi partitions, and the different demons algorithms. The additive demons algorithm provides better, or at least similar, reconstruction to the other demons algorithms. Moreover, the normalization in the empirical wavelet filters has no impact on the reconstruction performance for the different demons algorithms.
Despite the accuracy of the mapping estimation obtained in Table 1, the related wavelet reconstruction can be corrupted when the estimated filter bank does not entirely cover the Fourier domain. Notably, the transition width $\tau=0.1$ leads to lower signal-to-noise ratios than the transition width $\tau=0.2$. Indeed, in case of an inaccurate estimate of a mapping, $\tau=0.2$ ensures a larger covering of the Fourier domain by the wavelet filters than $\tau=0.1$. In contrast, the transition width $\tau=0.3$ can lead to artifacts in a symmetric wavelet filter $\chi_{n}$ due to overlaps of paired wavelet filters $\widehat{\psi}_{n}$ and $\widehat{\psi}_{-n}$, as confirmed by lower signal-to-noise ratios for the disk band-pass transform using the Watershed partition and a diffeomorphic demons algorithm compared to the transition width $\tau=0.2$.
@@ 4.5 Application to texture segmentation
In this section, we illustrate the relevance of the proposed Empirical wavelet transform in the segmentation of scanning tunneling microscope images. Analyzing the structure and chemical properties of self-assembled monolayers of organic molecules is the core of many applications in nanoscience and nanotechnology [18, 11]. These properties result in variations in the textures that are crucial to identify. Recently, Empirical Curvelet Wavelet Transform (EWT-Curvelet), obtained by partitioning the Fourier domain in scales and angles [10], has been shown to provide relevant texture features for scanning tunneling microscope images of self-assembled monolayers [12, 2, 13]. In particular, the EWT-Curvelet based segmentation outperforms the state-of-the-art methods for this task [2]. Therefore, we propose a comparison of our approach with the latter.
We perform the texture segmentation according to the procedure of [15], which consists of the following steps. First, we extract the cartoon and textural parts $u$ and $v$, respectively, of an image $f$ by solving the cartoon–texture decomposition model [1], defined as follows:
$\underset{u\in F,v\in G}{\mathrm{minimize}}\quad\|\nabla u\|_{1}+\mu\|v\|_{G}\quad\textrm{s.t.}\;f=u+v,$ (8)
where $F$ is the set of functions from ${\mathbb{R}}^{2}$ to ${\mathbb{R}}$ and $G$ is the set of oscillating functions. Following numerical analyses conducted by [5], the parameter controlling the oscillation degree of the textural part is set to $\mu=N_{\rm pixels}/2$. This problem is solved numerically using the algorithm described in [9]. Next, we compute the local energy of the empirical wavelet transform ${\mathcal{E}}_{\chi}^{v}$ of the textural part $v$, defined as the local average of the wavelet coefficients ${\mathcal{E}}_{\chi}^{v}(.,n)$ for each frequency band $n$, i.e.,
$\mathcal{T}^{{\mathcal{E}}}_{n}(x)=\frac{1}{N_{W}}\sum_{y\in W_{x}}\left|{\mathcal{E}}_{\chi}^{v}(y,n)\right|^{2},$ (9)
where $W_{x}$ is a window centered at the pixel $x$ of size set to $N_{W}=19\times 19$, following the work of [15]. Finally, the segmentation is obtained by performing a pixel-wise k-means clustering with the cityblock distance on the local energy (9), for preselected numbers of clusters $k$.
Figure 6 shows partitions of the Fourier spectrum of a scanning tunneling microscope image obtained by performing the Curvelet, Watershed and Voronoi methods with a scale-space step-size $s_{0}=0.2$ on the logarithm of the Fourier spectrum. The Curvelet method results in a large number of sets $\Omega_{n}$, namely $393$, while the Watershed and Voronoi methods extract as many sets $\Omega_{n}$ as detected harmonic modes, $17$ in this case.
@@
Figure 7 shows scanning tunneling microscope images, their corresponding textural parts, and their segmentation for the EWT-Curvelet and the proposed disk and square band-pass empirical wavelet transforms with $\tau=0.2$ for both Voronoi and Watershed partitioning. The proposed empirical wavelet transforms provide accurate and comparable texture masks for the different images. In all cases, the segmentation performance with the proposed approach is better than with the EWT-Curvelet, as can be clearly seen in the last two rows of Figure 7.
@@ 5 Discussion
As shown by numerical experiments, the additive demons algorithm provides accurate continuous mappings suitable for the empirical wavelet transform. The results show that this approach has satisfactory performances for reconstruction and texture extraction. In contrast, diffeomorphic mapping estimates are harder to obtain and take longer. This suggests that the use of normalization proposed in previous work [19] should be avoided in practice.
Moreover, although the proposed algorithm applies to any homogeneous or separable wavelet kernel and any Fourier partition, the selection of the wavelet kernel’s support and the partitioning method is crucial to optimize the mapping estimation performance. The study conducted in the present work shows that a Voronoi partition and a wavelet kernel defined on a disk result in a better performance overall.
The tools introduced in this work are ready for application on real-world data. In particular, they have been shown to provide relevant texture features for scanning tunneling microscope images for the different studied wavelet kernels. In addition, the number of features resulting from the Voronoi or Watershed partitioning methods is reasonable compared to the one obtained by the Curvelet transform used in [2].
Future work will include a thorough study of the stop criterion of the additive demons algorithm. In addition, an in-depth study of texture segmentation will also be carried out on a large labeled dataset, including a quantitative assessment of the impacts of the Fourier partitioning on the segmentation performance.
@@ 6 Conclusions
This work proposed an efficient algorithm for the empirical wavelet transform from any wavelet kernel based on demon registration. Among the different meticulously compared demons algorithms, the additive one has been shown to provide accurate numerical empirical wavelet systems with good properties of reconstruction, as validated for suitable wavelet kernels. The relevance of this approach for texture feature extraction has also been highlighted.